OpenStudy (anonymous):

What conversion factors should be used to convert 18mi/hr to ft/sec?

OpenStudy (debbieg):

Welcome to Open Study! Remember, in a conversion factor you are really just multiplying by "1", but you do it so that the units you don't want cancel, and the units you DO want remain. So for example, from miles to ft we have 1 mile = 5280 ft. So the conversion factor is either \[\Large \frac{ 1mi }{ 5280ft }\] or \[\Large \frac{ 5280 ft }{ 1mile }\] Where you want to go FROM miles TO feet, so you need miles to cancel, and you want feet left in the num'r, so you have: \[\Large x\text{miles}\cdot \frac{ 5280 ft }{ 1mile }=x\cancel{\text{miles}}\cdot \frac{ 5280 ft }{ 1\cancel{mile} }=5280x~ft\]

OpenStudy (debbieg):

Can you apply the above to come up with the hours to seconds conversion? :)